# UAV traffic scene understanding: A cross-spectral guided approach and a unified benchmark

### 11 Mar 2026 v1

@@ Abstract
Traffic scene understanding from unmanned aerial vehicle (UAV) platforms is crucial for intelligent transportation systems due to its flexible deployment and wide-area monitoring capabilities. However, existing methods face significant challenges in real-world surveillance, as their heavy reliance on optical imagery leads to severe performance degradation under adverse illumination conditions like nighttime and fog. Furthermore, current Visual Question Answering (VQA) models are restricted to elementary perception tasks, lacking the domain-specific regulatory knowledge required to assess complex traffic behaviors. To address these limitations, we propose a novel Cross-spectral Traffic Cognition Network (CTCNet) for robust UAV traffic scene understanding. Specifically, we design a Prototype-Guided Knowledge Embedding (PGKE) module that leverages high-level semantic prototypes from an external Traffic Regulation Memory (TRM) to anchor domain-specific knowledge into visual representations, enabling the model to comprehend complex behaviors and distinguish fine-grained traffic violations. Moreover, we develop a Quality-Aware Spectral Compensation (QASC) module that exploits the complementary characteristics of optical and thermal modalities to perform bidirectional context exchange, effectively compensating for degraded features to ensure robust representation in complex environments. In addition, we construct Traffic-VQA, the first large-scale optical-thermal infrared benchmark for cognitive UAV traffic understanding, comprising 8,180 aligned image pairs and 1.3 million question-answer pairs across 31 diverse types. Extensive experiments demonstrate that CTCNet significantly outperforms state-of-the-art methods in both cognition and perception scenarios. The dataset is available at https://github.com/YuZhang-2004/UAV-traffic-scene-understanding.
@@ 1 Introduction
Figure 1: Current challenges in UAV-based traffic VQA. (a) Data Gap. Existing datasets (top) rely on single-modal optical imagery for elementary perception, whereas practical surveillance (bottom) demands aligned OPT-TIR data for complex cognitive understanding. (b) Methodological Bottlenecks. General MLLMs struggle with the Domain Knowledge Gap, failing to interpret specific traffic rules (e.g., missing the "illegal" attribute of a U-turn), and the Multi-Modal Fusion Gap, where static fusion allows degraded optical noise to corrupt robust thermal features under adverse conditions. (c) Our Solution. The proposed CTCNet systematically bridges these gaps through the Traffic-VQA dataset, the PGKE module, and the QASC module.
@@
However, deploying UAV-VQA in real-world traffic surveillance presents significant scientific challenges. Beyond controlled ideal laboratory settings, practical UAV imagery is captured under unconstrained open-world conditions. These scenes are characterized by extreme viewpoint variations, dense distributions of tiny objects (e.g., miniaturized vehicles in wide-angle views), and severe vulnerability to environmental degradation such as nighttime darkness, glare, or dense fog [gujuanjuan, text-guided, Guidance_disentanglement]. Furthermore, interpreting complex traffic behaviors requires domain-specific regulatory knowledge, exposing a critical cognitive deficiency in current visual systems. As illustrated in Fig. 1, these factors lead to substantial visual feature corruption and semantic ambiguity, severely bottlenecking the perception capabilities of traditional visual models and limiting the practical deployment of UAV-VQA systems.
While existing VQA methods and Multimodal Large Language Models (MLLMs) designed for aerial platforms have driven significant progress in generalized spatial comprehension, they exhibit notable limitations. Representative methods such as GeoChat [geochat] and EarthGPT [earthgpt, EarthGPT-X] leverage pre-trained large models as unified interfaces to handle diverse interpretation tasks. However, a major vulnerability of these approaches is their predominant reliance on high-quality, single-modality optical (OPT) imagery. Since extracted feature representations are highly sensitive to input quality, optical signals are easily disrupted by environmental noise [Data_fusion]. Consequently, these models often suffer severe perception failures when applied to challenging, all-weather traffic scenarios where the optical signal is significantly degraded.
To mitigate the limitations of optical sensors, thermal infrared (TIR) imagery provides a complementary modality that captures heat signatures independent of ambient illumination, enabling effective performance in darkness and fog [gujuanjuan, Guidance_disentanglement, liu2024cross, zhao2025reflectance]. Fusing optical and TIR data represents a highly promising strategy for robust, all-weather perception. Mainstream multi-modal fusion strategies have evolved from simple static concatenation to sophisticated dynamic attention and correlation-driven mechanisms. Despite their success in low-level perception tasks such as object detection, these advanced cross-spectral fusion techniques have rarely been integrated into the high-level cognitive context of UAV-VQA, leaving a significant gap in robust multimodal feature representation for traffic behavior understanding.
@@
To address these combined challenges of multi-spectral data fusion and domain-specific cognitive depth, we propose a novel Cross-spectral Traffic Cognition Network (CTCNet). We decouple the problem into two complementary objectives: enhancing low-level perceptual robustness and injecting high-level cognitive context. Specifically, we design a Prototype-Guided Knowledge Embedding (PGKE) module that leverages an external Traffic Regulation Memory (TRM) constructed from expert knowledge. By retrieving and aligning high-level semantic prototypes with current visual features, the PGKE module injects domain-specific cognitive capabilities into the network. Simultaneously, we develop a Quality-Aware Spectral Compensation (QASC) module that orchestrates a bidirectional context exchange between optical and thermal modalities via a dynamic attention mechanism, allowing the network to selectively transfer discriminative features from the reliable modality to compensate for the degraded one, ensuring robust representation in complex environments. Beyond these methodological contributions, the research field also lacks a large-scale benchmark dataset with aligned multi-spectral imagery and cognitive annotations, which is crucial for training and evaluating robust traffic understanding models. To fill this gap, we construct Traffic-VQA, a large-scale OPT-TIR benchmark for UAV traffic understanding. It contains 8,180 meticulously aligned image pairs with over 1.3 million question-answer pairs, covering diverse environmental conditions (sunny, night, fog) and a comprehensive taxonomy of 31 question types, spanning from basic perception to complex violation understanding. Extensive experiments conducted on the proposed dataset demonstrate the superiority and effectiveness of CTCNet compared to existing state-of-the-art methods. The main contributions of this work are summarized as follows:
@@
  • Recognizing the vulnerability of single optical signals and the lack of domain-specific cognitive depth in current UAV traffic scene understanding, we propose the CTCNet framework to effectively harness complementary optical-thermal features and integrate domain-specific regulatory knowledge.
@@
  • Extensive experiments on the Traffic-VQA dataset validate the effectiveness of the proposed CTCNet. Compared with existing leading open-source MLLMs (e.g., Qwen2.5-VL [qwen2.5-vl], GeoChat [geochat]) and closed-source commercial models (e.g., GPT-4o [gpt4o]), our method successfully anchors domain-specific regulatory knowledge, achieving substantial improvements in comprehension accuracy for complex traffic behaviors across all-weather scenarios.
@@ 2 Related Work
The proposed CTCNet framework builds upon and advances three interrelated research areas: visual question answering, multimodal integration, and the cognitive capabilities of large language models. In what follows, we review representative works in each area and discuss their limitations in the context of UAV traffic understanding.
@@ 2.1 Visual Question Answering
2.2 Multimodal VQA
Multimodal VQA aims to leverage the complementary physics of heterogeneous sensors—such as optical and thermal or synthetic aperture radar (SAR) configurations—to increase informational density and analytical precision across diverse environmental contexts [earthgpt, soni2025earthdial]. Within this area, researchers have systematically addressed representational discrepancies originating from distinct imaging modalities and environmental degradation. To mitigate basic modality disparities, Li et al. [MFFENet] introduced a multiscale feature fusion and enhancement network designed to amplify foreground semantic objects, thereby improving the parsing of urban road networks under suboptimal illumination. To address the intrinsic modality gap, Zhou et al. [M-SpecGene] developed M-SpecGene, a generalized foundation model that uses a cross-modality structural sparsity metric to quantify information density and extract modality-invariant representations. Concurrently, Zhao et al. [CDDFuse] proposed CDDFuse, a correlation-driven architecture that decomposes features into modality-shared and modality-specific sub-components to improve cross-modal consistency.
Moving beyond static integration approaches, Zhang et al. [M2FNet] formulated M2FNet, which dynamically aggregates multi-spectral features via union-modal and cross-modal attention mechanisms to ensure robust object detection regardless of illumination variance. Furthermore, Xu et al. [U2Fusion] proposed a unified unsupervised framework, U2Fusion, which autonomously estimates the informational saliency of source imagery to guide adaptive feature preservation. Generative and prompt-guided strategies have also shown promise in bridging representational gaps. Advanced architectures such as AT-GAN [AT-GAN] and denoising diffusion models [DDFM] have been adapted to synthesize high-fidelity fused imagery while preserving intricate structural details. Specifically targeting unconstrained aerial environments, Yang et al. [GDNet] introduced GDNet for the disentanglement of optical guidance features, while PromptFusion [PromptFusion] harmonized multi-modality images guided by explicit semantic prompts through bi-level optimization.
Despite these promising results across various applications, the potential of advanced integration techniques within explicit cognitive contexts remains largely unexplored. Most existing multimodal methods rely on rigid integration strategies that lack the dynamic, non-destructive context exchange required to handle the extreme variance characteristic of all-weather traffic scenarios.
2.3 Multimodal Large Language Models
@@ 3 Methodology
In this section, we present the proposed CTCNet framework, a cognitive prototype-anchored architecture designed to augment MLLMs for robust traffic reasoning from UAV imagery. As illustrated in Fig. 2, our approach addresses two fundamental challenges in RSVQA: the domain knowledge gap inherent in interpreting specialized traffic behaviors, and the robust fusion of heterogeneous sensor modalities (OPT and TIR).
@@ 3.1 Overall Architecture
The overall architecture of CTCNet is illustrated in Fig. 2. The framework consists of a frozen MLLM backbone coupled with two parallel, task-specific enhancement branches: the PGKE module and the QASC module.
Figure 2: Overall framework of CTCNet for multi-spectral UAV traffic VQA. The architecture adopts a Gated Parallel Residual paradigm in which the frozen, pre-trained MLLM visual features are adaptively augmented by domain-specific residual knowledge generated by the PGKE and QASC modules. The learnable gating parameters $\alpha$ and $\beta$ regulate the intensity of cognitive and multimodal context injection.
Given a well-aligned pair of optical and TIR images, denoted as $I_{opt}\in\mathbb{R}^{H\times W\times 3}$ and $I_{th}\in\mathbb{R}^{H\times W\times 3}$, alongside a natural language query $Q$, CTCNet aims to generate a comprehensive textual answer $A$. To harness the generalized inferential capabilities of large-scale pre-training while adapting to the specificities of UAV traffic environments, we employ a frozen MLLM backbone (e.g., Qwen-VL) augmented with a non-invasive residual enhancement mechanism.
@@ 3.5 Loss Function
Following standard MLLM training protocols, CTCNet is optimized via an auto-regressive language modeling objective. Given the sequence of ground-truth answer tokens $\mathbf{y}=\{y_{1},y_{2},...,y_{T}\}$, the training objective is to minimize the negative log-likelihood:
@@ 4.1 Review of Existing Datasets
Cognition-Oriented and Large-Scale Benchmarks. Recent efforts have aimed to align datasets with the capabilities of MLLMs. CRSVQA [MQVQA] introduces complex, multi-tiered questions to benchmark advanced question-driven systems. RSGPT [rsgpt] establishes a benchmark for simultaneous image captioning and VQA. EarthVQA [earthvqa] emphasizes complex relational analysis for urban planning applications, whereas LRS-VQA [When_Large] targets gigapixel-level ultra-large imagery interpretation. EarthGPT [earthgpt] integrates a diverse set of tasks but still relies fundamentally on optical data. Critically, none of these benchmarks provide well-aligned multimodal (optical and TIR) data for robust all-weather perception, nor do they target the fine-grained cognitive tasks (e.g., traffic violation detection, vehicle behavior analysis) that are central to intelligent transportation systems.
@@
Dataset Num Img Num QA Num Qst Type Main QA Generation Incl. Cognitive Modal Key Tasks RSVQA-LR [RSVQA] 772 77k 5 Template OPT Counting, Presence RSVQA-HR [RSVQA] 10,659 1,066k 5 Template OPT Counting, Presence VQA-TextRS [VQA-TextRS] 2,144 6.2k 4 Manual OPT Object Recognition RSIVQA [rsivqa] 37,264 111k 9 Template OPT Object Recognition CRSVQA [MQVQA] 4,639 4.6k 3 Manual OPT Multistep Analysis FloodNet-VQA [floodnet] 2,188 7.4k 4 Manual OPT Disaster Assessment SAM-VQA [SAM-VQA] 2,348 10.5k 7 Template OPT Damage Analysis CDVQA [CDVQA] 4,662 >>122k 5 Template OPT Change Detection RSIEval [rsgpt] 100 0.9k 4 Manual OPT Captioning & VQA EarthVQA [earthvqa] 6,000 209k 3 Manual + Template OPT Relational Analysis OSVQA [text-guided] 6,008 pairs 1,037k 16 Manual + Template OPT & SAR Multimodal Perception Traffic-VQA (Ours) 8,180 pairs 1,301k 31 Manual + LLM OPT & TIR Traffic Cognition
Figure 7: Statistical distribution of the Traffic-VQA dataset. Left: Breakdown of the 8,180 image pairs by illumination conditions, weather, and scene type. Right: Distribution of the 31 question types, highlighting the significant proportion of high-level cognitive tasks (20.7%) and specialized multimodal queries.

### 14 Mar 2026 v2

# UAV traffic scene understanding: A regulation embedded multi-modal network and a unified benchmark

@@ Abstract
Traffic scene understanding from unmanned aerial vehicle (UAV) platforms is crucial for intelligent transportation systems due to its flexible deployment and wide-area monitoring capabilities. However, existing methods face significant challenges in real-world surveillance, as their heavy reliance on optical imagery leads to severe performance degradation under adverse illumination conditions like nighttime and fog. Furthermore, current Visual Question Answering (VQA) models are restricted to elementary perception tasks, lacking the domain-specific regulatory knowledge required to assess complex traffic behaviors. To address these limitations, we propose a novel Multi-modal Traffic Cognition Network (MTCNet) for robust UAV traffic scene understanding. Specifically, we design a Prototype-Guided Knowledge Embedding (PGKE) module that leverages high-level semantic prototypes from an external Traffic Regulation Memory (TRM) to anchor domain-specific knowledge into visual representations, enabling the model to comprehend complex behaviors and distinguish fine-grained traffic violations. Moreover, we develop a Quality-Aware Spectral Compensation (QASC) module that exploits the complementary characteristics of optical and thermal modalities to perform bidirectional context exchange, effectively compensating for degraded features to ensure robust representation in complex environments. In addition, we construct Traffic-VQA, the first large-scale optical-thermal infrared benchmark for cognitive UAV traffic understanding, comprising 8,180 aligned image pairs and 1.3 million question-answer pairs across 31 diverse types. Extensive experiments demonstrate that MTCNet significantly outperforms state-of-the-art methods in both cognition and perception scenarios. The dataset is available at https://github.com/YuZhang-2004/UAV-traffic-scene-understanding.
@@ 1 Introduction
Figure 1: Current challenges in UAV-based traffic VQA. (a) Data Gap. Existing datasets (top) rely on single-modal optical imagery for elementary perception, whereas practical surveillance (bottom) demands aligned OPT-TIR data for complex cognitive understanding. (b) Methodological Bottlenecks. General MLLMs struggle with the Domain Knowledge Gap, failing to interpret specific traffic rules (e.g., missing the "illegal" attribute of a U-turn), and the Multi-Modal Fusion Gap, where static fusion allows degraded optical noise to corrupt robust thermal features under adverse conditions. (c) Our Solution. The proposed MTCNet systematically bridges these gaps through the Traffic-VQA dataset, the PGKE module, and the QASC module.
@@
However, deploying UAV-VQA in real-world traffic surveillance presents significant scientific challenges. Beyond controlled ideal laboratory settings, practical UAV imagery is captured under unconstrained open-world conditions. These scenes are characterized by extreme viewpoint variations, dense distributions of tiny objects (e.g., miniaturized vehicles in wide-angle views), and severe vulnerability to environmental degradation such as nighttime darkness, glare, or dense fog [Guidance_disentanglement]. Furthermore, interpreting complex traffic behaviors requires domain-specific regulatory knowledge, exposing a critical cognitive deficiency in current visual systems. As illustrated in Fig. 1, these factors lead to substantial visual feature corruption and semantic ambiguity, severely bottlenecking the perception capabilities of traditional visual models and limiting the practical deployment of UAV-VQA systems.
While existing VQA methods and Multi-modal Large Language Models (MLLMs) designed for aerial platforms have driven significant progress in generalized spatial comprehension, they exhibit notable limitations. Representative methods such as GeoChat [geochat] and EarthGPT [earthgpt, EarthGPT-X] leverage pre-trained large models as unified interfaces to handle diverse interpretation tasks. However, a major vulnerability of these approaches is their predominant reliance on high-quality, single-modality optical (OPT) imagery. Since extracted feature representations are highly sensitive to input quality, optical signals are easily disrupted by environmental noise [Data_fusion]. Consequently, these models often suffer severe perception failures when applied to challenging, all-weather traffic scenarios where the optical signal is significantly degraded.
To mitigate the limitations of optical sensors, thermal infrared (TIR) imagery provides a complementary modality that captures heat signatures independent of ambient illumination, enabling effective performance in darkness and fog [gujuanjuan, Guidance_disentanglement, liu2024cross, zhao2025reflectance]. Fusing optical and TIR data represents a highly promising strategy for robust, all-weather perception. Mainstream multi-modal fusion strategies have evolved from simple static concatenation to sophisticated dynamic attention and correlation-driven mechanisms. Despite their success in low-level perception tasks such as object detection, these advanced multi-modal fusion techniques have rarely been integrated into the high-level cognitive context of UAV-VQA, leaving a significant gap in robust multi-modal feature representation for traffic behavior understanding.
@@
To address these combined challenges of multi-spectral data fusion and domain-specific cognitive depth, we propose a novel Multi-modal Traffic Cognition Network (MTCNet). We decouple the problem into two complementary objectives: enhancing low-level perceptual robustness and injecting high-level cognitive context. Specifically, we design a Prototype-Guided Knowledge Embedding (PGKE) module that leverages an external Traffic Regulation Memory (TRM) constructed from expert knowledge. By retrieving and aligning high-level semantic prototypes with current visual features, the PGKE module injects domain-specific cognitive capabilities into the network. Simultaneously, we develop a Quality-Aware Spectral Compensation (QASC) module that orchestrates a bidirectional context exchange between optical and thermal modalities via a dynamic attention mechanism, allowing the network to selectively transfer discriminative features from the reliable modality to compensate for the degraded one, ensuring robust representation in complex environments. Beyond these methodological contributions, the research field also lacks a large-scale benchmark dataset with aligned multi-spectral imagery and cognitive annotations, which is crucial for training and evaluating robust traffic understanding models. To fill this gap, we construct Traffic-VQA, a large-scale OPT-TIR benchmark for UAV traffic understanding. It contains 8,180 meticulously aligned image pairs with over 1.3 million question-answer pairs, covering diverse environmental conditions (sunny, night, fog) and a comprehensive taxonomy of 31 question types, spanning from basic perception to complex violation understanding. Extensive experiments conducted on the proposed dataset demonstrate the superiority and effectiveness of MTCNet compared to existing state-of-the-art methods. The main contributions of this work are summarized as follows:
@@
  • Recognizing the vulnerability of single optical signals and the lack of domain-specific cognitive depth in current UAV traffic scene understanding, we propose the MTCNet framework to effectively harness complementary optical-thermal features and integrate domain-specific regulatory knowledge.
@@
  • Extensive experiments on the Traffic-VQA dataset validate the effectiveness of the proposed MTCNet. Compared with existing leading open-source MLLMs (e.g., Qwen2.5-VL [qwen2.5-vl], GeoChat [geochat]) and closed-source commercial models (e.g., GPT-4o [gpt4o]), our method successfully anchors domain-specific regulatory knowledge, achieving substantial improvements in comprehension accuracy for complex traffic behaviors across all-weather scenarios.
@@ 2 Related Work
The proposed MTCNet framework builds upon and advances three interrelated research areas: visual question answering, multi-modal integration, and the cognitive capabilities of large language models. In what follows, we review representative works in each area and discuss their limitations in the context of UAV traffic understanding.
@@ 2.1 Visual Question Answering
2.2 Multi-modal VQA
Multi-modal VQA aims to leverage the complementary physics of heterogeneous sensors—such as optical and thermal or synthetic aperture radar (SAR) configurations—to increase informational density and analytical precision across diverse environmental contexts [earthgpt, soni2025earthdial]. Within this area, researchers have systematically addressed representational discrepancies originating from distinct imaging modalities and environmental degradation. To mitigate basic modality disparities, Li et al. [MFFENet] introduced a multiscale feature fusion and enhancement network designed to amplify foreground semantic objects, thereby improving the parsing of urban road networks under suboptimal illumination. To address the intrinsic modality gap, Zhou et al. [M-SpecGene] developed M-SpecGene, a generalized foundation model that uses a cross-modality structural sparsity metric to quantify information density and extract modality-invariant representations. Concurrently, Zhao et al. [CDDFuse] proposed CDDFuse, a correlation-driven architecture that decomposes features into modality-shared and modality-specific sub-components to improve cross-modal consistency.
Moving beyond static integration approaches, Zhang et al. [M2FNet] formulated M2FNet, which dynamically aggregates multi-spectral features via union-modal and cross-modal attention mechanisms to ensure robust object detection regardless of illumination variance. Furthermore, Xu et al. [U2Fusion] proposed a unified unsupervised framework, U2Fusion, which autonomously estimates the informational saliency of source imagery to guide adaptive feature preservation. Generative and prompt-guided strategies have also shown promise in bridging representational gaps. Advanced architectures such as AT-GAN [AT-GAN] and denoising diffusion models [DDFM] have been adapted to synthesize high-fidelity fused imagery while preserving intricate structural details. Specifically targeting unconstrained aerial environments, Zhao et al. [Guidance_disentanglement] introduced GDNet for the disentanglement of optical guidance features, while PromptFusion [PromptFusion] harmonized multi-modality images guided by explicit semantic prompts through bi-level optimization.
Despite these promising results across various applications, the potential of advanced integration techniques within explicit cognitive contexts remains largely unexplored. Most existing multi-modal methods rely on rigid integration strategies that lack the dynamic, non-destructive context exchange required to handle the extreme variance characteristic of all-weather traffic scenarios.
2.3 Multi-modal Large Language Models
@@ 3 Methodology
In this section, we present the proposed MTCNet framework, a cognitive prototype-anchored architecture designed to augment MLLMs for robust traffic reasoning from UAV imagery. As illustrated in Fig. 2, our approach addresses two fundamental challenges in RSVQA: the domain knowledge gap inherent in interpreting specialized traffic behaviors, and the robust fusion of heterogeneous sensor modalities (OPT and TIR).
@@ 3.1 Overall Architecture
The overall architecture of MTCNet is illustrated in Fig. 2. The framework consists of a frozen MLLM backbone coupled with two parallel, task-specific enhancement branches: the PGKE module and the QASC module.
Figure 2: Overall framework of MTCNet for multi-spectral UAV traffic VQA. The architecture adopts a Gated Parallel Residual paradigm in which the frozen, pre-trained MLLM visual features are adaptively augmented by domain-specific residual knowledge generated by the PGKE and QASC modules. The learnable gating parameters $\alpha$ and $\beta$ regulate the intensity of cognitive and multi-modal context injection.
Given a well-aligned pair of optical and TIR images, denoted as $I_{opt}\in\mathbb{R}^{H\times W\times 3}$ and $I_{th}\in\mathbb{R}^{H\times W\times 3}$, alongside a natural language query $Q$, MTCNet aims to generate a comprehensive textual answer $A$. To harness the generalized inferential capabilities of large-scale pre-training while adapting to the specificities of UAV traffic environments, we employ a frozen MLLM backbone (e.g., Qwen-VL) augmented with a non-invasive residual enhancement mechanism.
@@ 3.5 Loss Function
Following standard MLLM training protocols, MTCNet is optimized via an auto-regressive language modeling objective. Given the sequence of ground-truth answer tokens $\mathbf{y}=\{y_{1},y_{2},...,y_{T}\}$, the training objective is to minimize the negative log-likelihood:
@@ 4.1 Review of Existing Datasets
Cognition-Oriented and Large-Scale Benchmarks. Recent efforts have aimed to align datasets with the capabilities of MLLMs. CRSVQA [MQVQA] introduces complex, multi-tiered questions to benchmark advanced question-driven systems. RSGPT [rsgpt] establishes a benchmark for simultaneous image captioning and VQA. EarthVQA [earthvqa] emphasizes complex relational analysis for urban planning applications, whereas LRS-VQA [When_Large] targets gigapixel-level ultra-large imagery interpretation. EarthGPT [earthgpt] integrates a diverse set of tasks but still relies fundamentally on optical data. Critically, none of these benchmarks provide well-aligned multi-modal (optical and TIR) data for robust all-weather perception, nor do they target the fine-grained cognitive tasks (e.g., traffic violation detection, vehicle behavior analysis) that are central to intelligent transportation systems.
@@
Dataset Num Img Num QA Num Qst Type Main QA Generation Incl. Cognitive Modal Key Tasks RSVQA-LR [RSVQA] 772 77k 5 Template OPT Counting, Presence RSVQA-HR [RSVQA] 10,659 1,066k 5 Template OPT Counting, Presence VQA-TextRS [VQA-TextRS] 2,144 6.2k 4 Manual OPT Object Recognition RSIVQA [rsivqa] 37,264 111k 9 Template OPT Object Recognition CRSVQA [MQVQA] 4,639 4.6k 3 Manual OPT Multistep Analysis FloodNet-VQA [floodnet] 2,188 7.4k 4 Manual OPT Disaster Assessment SAM-VQA [SAM-VQA] 2,348 10.5k 7 Template OPT Damage Analysis CDVQA [CDVQA] 4,662 >>122k 5 Template OPT Change Detection RSIEval [rsgpt] 100 0.9k 4 Manual OPT Captioning & VQA EarthVQA [earthvqa] 6,000 209k 3 Manual + Template OPT Relational Analysis OSVQA [text-guided] 6,008 pairs 1,037k 16 Manual + Template OPT & SAR Multi-modal Perception Traffic-VQA (Ours) 8,180 pairs 1,301k 31 Manual + LLM OPT & TIR Traffic Cognition
Figure 7: Statistical distribution of the Traffic-VQA dataset. Left: Breakdown of the 8,180 image pairs by illumination conditions, weather, and scene type. Right: Distribution of the 31 question types, highlighting the significant proportion of high-level cognitive tasks (20.7%) and specialized multi-modal queries.